OpenStudy (sasogeek):

if i'm supposed to write a given log as a single logarithm, would this be right given \(\large (x+6)log_4 3\) ans. \(\large log_4 3^x + log_4 3^6 \) ?

OpenStudy (across):

There's one last step.

OpenStudy (sasogeek):

\(\large xlog_4 3+6log_4 3 \) ?

OpenStudy (amistre64):

a shorter route may be:\[n~log(a)=log(a^n)\]

OpenStudy (across):

Recall that \(\log a+\log b=\log ab\).

mathslover (mathslover):